What is the dissociation reaction for this solution?

Consider a 0.35 M pyridine solution, C5H5N(aq), that has a pH of 9.38.

  1. What is the dissociation reaction for this solution? hint: apply rules for bronsted-lowry theory
  2. What is the Kb for C5H5N(aq)?
Answer #1

1)
C5H5N (aq) + H2O (l)     <->     C5H5NH+(aq)   +   OH-(aq)


2)
use:
pH = -log [H+]
9.38 = -log [H+]
[H+] = 4.169*10^-10 M

use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(4.169*10^-10)
[OH-] = 2.399*10^-5 M
C5H5N dissociates as:

C5H5N +H2O     ----->     C5H5NH+   +   OH-
0.35                   0         0
0.35-x                 x         x


Kb = [C5H5NH+][OH-]/[C5H5N]
Kb = x*x/(c-x)
Kb = 2.399*10^-5*2.399*10^-5/(0.35-2.399*10^-5)
Kb = 1.644*10^-9
Answer: 1.6*10^-9